a person consumes a cup of rice pudding containing 50g of rice 200g of milk and 50g of sugar calculate the corresponding energy value knowing that 1g of carbohydrate provides 3.75 Kcal 1g gram of fat provides 9Kcal and 1g of priteine ​​provides 4Kcal
1
Expert's answer
2021-03-22T06:26:14-0400

Total amount consumed is 50+ 200+50= 300g

Energy from rice= 187.5kcal

Energy from milk=200× 9= 1800kcal

Energy from sugar is 200kcal

Total energy= 2187.5kcal